# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ 7. Numerical Experiments
We compare the optimal bounds for specific local randomizers under the general clone paradigm with existing bounds. The baselines include two bounds from the privacy blanket framework, derived using Hoeffding’s and Bennett’s inequalities, respectively (Balle et al., 2019), as well as the numerical bounds from the standard clone paradigm (Feldman et al., 2022). We utilized the open-source implementations released by the respective papers.
We present results for four commonly used local randomizers: $k$-Randomized Response (Warner, 1965) ($k=10$), Binary Local Hash (Wang et al., 2017), RAPPOR (Erlingsson et al., 2014), and Optimized Unary Encoding (OUE) (Wang et al., 2017). The experimental results are shown in Figure 2 and Figure 3. The lower bounds are computed using the method described in Section 5.3. In the legends of the plots, “$k$-joint” refers to the joint composition of $k$ local randomizers, each satisfying $\frac{\epsilon_{0}}{k}$-LDP. In all experiments, the discretization interval length $l$ of our algorithm is set as $\frac{e^{\epsilon_{0}}-1}{1200}$.
The experimental results show that our upper bounds consistently outperform existing methods. Furthermore, the gap between our upper and lower bounds is generally small, indicating the tightness and reliability of the computed results. This also validates that our choice of discretization interval length $l$ offers sufficient precision for practical use.
In addition to its accuracy, our algorithm is highly efficient: it generates a full curve in approximately 30 seconds, whereas the numerical algorithm used for the standard clone requires around 5 minutes. It is worth noting that the numerical generic bound of the standard clone is weaker than the specific one provided by the privacy blanket using Bennett’s inequality in our experiments.
@@
Our results show that computing bounds specific to joint compositions leads to significantly tighter amplification bounds compared to generic methods. This highlights the importance and advantage of our algorithm in accurately analyzing privacy in practical multi-attribute settings. Furthermore, we observe that, under a fixed total privacy budget $\epsilon_{0}$, the privacy amplification effect achieved through shuffling becomes stronger as the number of composed randomizers $k$ increases.
10-RR, low $\epsilon_{0}$
10-RR, large $\epsilon_{0}$
BLH, low $\epsilon_{0}$
BLH, large $\epsilon_{0}$
Figure 2. Experimental Results: RR and BLH
RAPPOR, low $\epsilon_{0}$
RAPPOR, large $\epsilon_{0}$
OUE, low $\epsilon_{0}$
OUE, large $\epsilon_{0}$
Figure 3. Experimental Results: RAPPOR and OUE
8. Discussion: Beyond the General Clone
@@
A promising direction is to identify the most vulnerable neighboring dataset pair $(X^{0},X^{1})$ such that $D_{e^{\epsilon}}\left(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\right)$ is maximized among all neighboring pairs of size $n$. If, for a local randomizer $\mathcal{R}$, one can prove that a specific pair $(X^{0}_{v},X^{1}_{v})$ consistently maximizes the divergence for every $\epsilon$, then this would yield the exact privacy amplification bound for $\mathcal{R}$ under shuffling.
@@
9. Conclusion

### v5

@@ 6 Numerical Experiments
We compare our computed bounds for several common local randomizers against existing bounds from prior work. The baselines include two bounds derived from the privacy blanket framework using Hoeffding’s and Bennett’s inequalities, respectively [BBGN19], as well as numerical bounds from the standard clone paradigm [FMT22]. We use the publicly available implementations released by the respective authors.
Experimental setup.
We present results for five widely used local randomizers: $k$-ary randomized response (with $k=10$[War65], Binary Local Hash (BLH) [WBLJ17], RAPPOR [EPK14], Optimized Unary Encoding (OUE) [WBLJ17], and the Laplace mechanism over $\{0,1\}$ [Dwo06]. For simplicity, we use the asymptotic forms of GPARV for BLH, RAPPOR, and OUE in the regime where the domain size $D\gg 1$. As an illustrative example, we also report how the upper bound for RAPPOR changes with different values of $D$. As shown in Fig. 5, when $D\geq 5$, the upper bound curves become nearly indistinguishable. Furthermore, for $D\geq 3$, where $x_{2}\in\mathbb{X}\setminus\{x_{1}^{0},x_{1}^{1}\}$ can be selected, the lower-bound GPARV for the above local randomizers becomes independent of $D$.
Results and observations.
The main results are presented in Fig. 2 and Fig. 3. Lower bounds are computed as described in Section 5.3. In the legends, “$k$-joint” denotes the joint composition of $k$ local randomizers, each satisfying $\frac{\varepsilon_{0}}{k}$-LDP. The discretization interval $l$ used in our FFT-based algorithm is set to $\frac{e^{\varepsilon_{0}}-1}{1200}$ when $\varepsilon_{0}=0.1$, and $\frac{e^{\varepsilon_{0}}-1}{1000}$ when $\varepsilon_{0}=4.0$.
Our results show that the computed upper bounds consistently outperform all baselines. Moreover, the gap between our upper and lower bounds is generally small, confirming the tightness and reliability of our analysis. These results also validate that the chosen discretization step $l$ provides sufficient numerical precision in practice. In addition to accuracy, our algorithm is highly efficient: a full amplification curve is generated in approximately 30 seconds, compared to roughly 5 minutes required by the standard clone’s numerical method.
Impact of joint composition.
@@
Our findings further highlight the advantage of computing specific bounds for joint compositions. The resulting amplification is significantly tighter than those derived from generic methods, illustrating the strength of our approach in multi-attribute settings. Moreover, under a fixed total privacy budget $\varepsilon_{0}$, we observe that the amplification effect improves as the number of composed randomizers $k$ increases.
Impact of parallel composition.
We evaluate the parallel composition of 10-ary randomized response (10-RR) and BLH, where each mechanism is selected with equal probability $\boldsymbol{p}=[0.5,0.5]$. As shown in Fig. 4, the amplification curve of the parallel composition lies between those of the individual mechanisms, aligning well with theoretical expectations.
Poisson subsampling in the shuffle model.
As an illustrative case, Fig. 6 presents the upper bound computed by our method when applied to Poisson subsampling in the shuffle model, instantiated with 10-RR and analyzed via parallel composition. Notably, the corresponding lower bound closely matches the upper bound, demonstrating the tightness of our analysis in this setting as well.
7 Discussion: Beyond the General Clone
@@
A promising direction is to identify the most vulnerable neighboring dataset pair $(X^{0},X^{1})$ such that $D_{e^{\varepsilon}}\left(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\right)$ is maximized among all neighboring pairs of size $n$. If, for a local randomizer $\mathcal{R}$, one can prove that a specific pair $(X^{0}_{v},X^{1}_{v})$ consistently maximizes the divergence for every $\varepsilon$, then this would yield the exact privacy amplification bound for $\mathcal{R}$ under shuffling.
@@
8 Limitation
When a local randomizer satisfies input symmetry, it is sufficient to evaluate $G_{\varepsilon}^{x,x^{\prime}}$ for any fixed neighboring pair $(x,x^{\prime})$; the resulting privacy bound then applies to all neighboring datasets. However, in the absence of input symmetry, one must, in principle, evaluate $G_{\varepsilon}^{x,x^{\prime}}$ for all possible input pairs and take the maximum to ensure worst-case validity.
For the Laplace mechanism over $[0,1]$, we currently do not know how to prove or disprove whether the pair $(x=0,x^{\prime}=1)$ constitutes the worst case under the blanket decomposition. We leave this question as future work.
In the case of joint composition, a subtle issue arises regarding input symmetry. Even if each sub-local randomizer individually satisfies input symmetry, their joint composition introduces additional complications. Specifically, consider two neighboring inputs $(x^{1},x^{2},\dots,x^{m})$ and $(\bar{x}^{1},\bar{x}^{2},\dots,\bar{x}^{m})$ such that $x^{i}\neq\bar{x}^{i}$ for all $i$. In this case, the pair is symmetric, and it suffices to evaluate the GPARV on a single such instance.
However, when the adjacency relation is defined over the full product domain $\mathbb{X}^{1}\times\mathbb{X}^{2}\times\dots\times\mathbb{X}^{m}$, one must also consider neighboring pairs that differ in only a subset of coordinates. If all $\mathcal{R}_{i}$ are identical, we need to separately consider adjacency relations with Hamming distances $1,2,\dots,m$. For example, in the joint composition of three $\varepsilon_{0}/4$-DP $k$-RR mechanisms, we must evaluate the GPARV for the following input pairs:
$\big{(}(0,0,0,0),(1,1,1,1)\big{)},\quad\big{(}(0,0,0,0),(0,1,1,1)\big{)},\quad\big{(}(0,0,0,0),(0,0,1,1)\big{)},\quad\big{(}(0,0,0,0),(0,0,0,1)\big{)}.$ .
Among them, the pair $\big{(}(0,0,0,0),(1,1,1,1)\big{)}$ is theoretically expected to yield the GPARV with the highest variance, and our experimental results confirm that it produces the largest upper bound (see Fig. 7). Whether this observation can be formally proven remains an open question for future work.
9 Conclusion